A sample containing 2.80 grams of O2 gas has an initial volume of 15.0 mL What is the final volume when 0.500 moles of O2 is added at constant temperature and pressure?
1
Expert's answer
2019-02-26T06:20:58-0500

The equation of Ideal Gas Law PV = nRT shows that volume is directly proportional to amount of substance at constant temperature and pressure.

n= m/M

n(O2) = 2.80g/ 32.00 g/mol = 0.0875 mol

nfinal(O2) = 0.0875 mol + 0.500 mol = 0.5875 mol

Solve the proportion:

0.0875 mol of O2 occupy 15.0 mL

0.5875 mol occupy x mL

0.0875/0.5875 = 15.0/x

x= 100.7 mL

Answer: 100.7 mL